### v2

# Integrating Weather Station Data and Radar for Precipitation Nowcasting: SmaAt-fUsion and SmaAt-Krige-GNet

Abstract
Short-term precipitation nowcasting is essential for flood management, transportation, energy system operations, and emergency response. However, many existing models fail to fully exploit the extensive atmospheric information available, relying primarily on precipitation data alone. This study examines whether integrating multi-variable weather-station measurements with radar can enhance nowcasting skill and introduces two complementary architectures that integrate multi-variable station data with radar images. The SmaAt-fUsion model extends the SmaAt-UNet framework by incorporating weather station data through a convolutional layer, integrating it into the bottleneck of the network; The SmaAt-Krige-GNet model combines precipitation maps with weather station data processed using Kriging, a geo-statistical interpolation method, to generate variable-specific maps. These maps are then utilized in a dual-encoder architecture based on SmaAt-GNet, allowing multi-level data integration . Experimental evaluations were conducted using four years (2016–2019) of weather station and precipitation radar data from the Netherlands. Results demonstrate that SmaAt-Krige-GNet outperforms the standard SmaAt-UNet, which relies solely on precipitation radar data, in low precipitation scenarios, while SmaAt-fUsion surpasses SmaAt-UNet in both low and high precipitation scenarios. This highlights the potential of incorporating discrete weather station data to enhance the performance of deep learning-based weather nowcasting models.
keywords:
Weather station data , radar data , fusion , deep learning , precipitation nowcasting
journal: arXiv
1 Introduction
Precipitation nowcasting refers to the short-term forecasting of rainfall intensity, typically within a timeframe ranging from a few minutes to several hours. Accurate nowcasting is crucial for various weather-dependent sectors, including flood management, transportation, energy distribution, and emergency response li2024quantitative; wilson2010nowcasting. However, although numerical weather prediction (NWP) and radar-based methods have made notable progress, there remains improvement in operational settings. NWP models utilize physical atmospheric properties to generate multiple precipitation scenarios, but they are computationally expensive and often impractical for short-term forecasts due to their sensitivity to initial conditions and require high computations bauer2015quiet; sun2014use; wilks2011statistical. This gap poses a growing threat to the effectiveness of early warning systems, particularly in densely populated or infrastructure-critical regions such as the Netherlands. The effectiveness of nowcasting models depends on their ability to capture the complex, dynamic nature of precipitation patterns across different spatial and temporal scales.
Recent advancements in deep learning have led to the development of advanced data-driven models that utilize historical weather data to capture complex spatiotemporal patterns mehrkanoon2019deep; ren2021deep; vatamany2025graph; price2025probabilistic; ham2023anthropogenic; allen2025end. Convolutional Neural Networks (CNNs) based models, in particular, have demonstrated promising results in precipitation nowcasting by effectively extracting spatial features from radar images ayzel2019all; ayzel2020rainnet; jiang2024hybrid. However, these models almost exclusively rely on radar reflectivity data, overlooking other atmospheric drivers, such as temperature, humidity, pressure, and wind, which fundamentally influence precipitation formation. As a result, they are unable to capture the localized signals that often precede rainfall initiation. Consequently, radar-only systems may fail to capture pre-convective signals, limiting both accuracy and generalization under rapidly changing weather conditions.
Some recent studies have attempted to incorporate multiple atmospheric variables into their models to improve prediction robustness wang2025enhancing; cao2025enhancing. For instance, the authors in kaparakis2023wf introduced WF-UNet, a two-stream UNet architecture that processes both precipitation maps and wind speed maps to assess the impact of wind on precipitation nowcasting. However, such approaches typically depend on spatially continuous maps rather than integrating discrete station-based observations. Weather stations provide highly localized, multi-variable atmospheric observations that offer valuable insights into the physical processes behind rainfall. However, station data alone lacks spatial continuity, making it challenging to use directly in deep learning models designed for grid-based data. Our work aims to bridge this gap by learning from both discrete weather station data and radar-based precipitation maps.
Among various spatial interpolation techniques, Kriging Isaaks1989 has been widely used as a geostatistical method for generating spatially continuous fields from discrete observations. It estimates both spatial values and associated uncertainties through a variogram model 8102132; brown2022variogram. This makes Kriging performs well for smoothly varying meteorological variables and provides interpretable spatial correlations. However, it also suffers from several weaknesses when applied to high-resolution or rapidly evolving rainfall processes. It assumes spatial stationarity, leading to performance degradation during convective events; it tends to over-smooth local extremes, thereby reducing the representation of sharp rainfall gradients. These strengths and weaknesses highlight the need to combine Kriging interpolation with the deep learning to produce more realistic and dynamically consistent precipitation fields.
To address this limitation, this study explores the potential benefits of incorporating additional meteorological data into deep learning-based nowcasting models. Specifically, we introduce two novel architectures, i.e. SmaAt-fUsion and SmaAt-Krige-GNet, that integrate multi-variable weather station data alongside precipitation radar data. The key contributions of this work are as follows.
  1. 1.
    Integrate multi-source dataset: We integrate four years (2016–2019) of radar-based precipitation data with multi-variable observations from 22 KNMI weather stations across the Netherlands. The datasets were temporally aligned and standardized to enable pixel-level fusion between radar grids and discrete station data, provides a rich source of information for improving precipitation nowcasting.
  2. 2.
    Development of SmaAt-fUsion: This model extends the SmaAt-UNet framework trebing2021smaat by incorporating weather station data into the bottleneck of the SmaAt-UNet model, where high-level representations are formed.
  3. 3.
    Development of SmaAt-Krige-GNet: To explicitly capture spatial relationships between weather stations, we employ the Kriging algorithm, a geo-statistical interpolation method, to generate spatially continuous maps from the collected discrete weather station data. These Kriging-derived maps serve as an additional data input to SmaAt-Krige-GNet, which extends the SmaAt-UNet model by incorporating a dual-encoder architecture. The Kriging-based stream is concatenated with the precipitation stream at multiple stages of the network, ensuring that the model learns from both data sources at various levels of abstraction.
2 Related Work
Quantitative Precipitation Nowcasting (QPN) has experienced a fundamental transformation from statistical extrapolation and optical-flow approaches toward deep learning-based forecasting 10220887. Classical optical-flow systems such as RainyMotion ayzel2019optical and PySTEPS pulkkinen2019pysteps estimate the displacement of radar echoes between consecutive frames and have long served as operational baselines due to their simplicity and physical interpretability. Early deep learning architectures have been successfully applied in weather forecasting, such as Recurrent Neural Networks (RNNs) singh2019weather, Long Short-Term Memory (LSTM) networks zaytar2016sequence, Convolutional LSTM (ConvLSTM) shi2015convolutional, Convolutional Neural Networks (CNNs) ayzel2020rainnet, Graph Neural Networks (GNNs) simeunovic2021spatio, and encoder-decoder frameworks larraondo2019data. All of which have shown promising results in capturing complex spatio-temporal precipitation patterns.
Building upon these advances, the UNet architecture ronneberger2015unetconvolutionalnetworksbiomedical, originally developed for medical image segmentation, has demonstrated remarkable effectiveness in precipitation nowcasting due to its encoder–decoder structure with skip connections, which enables the preservation of fine-grained spatial details. Several variations of UNet have been proposed to improve both performance and computational efficiency for weather prediction tasks. For instance, SmaAt-UNet trebing2021smaat, an enhanced lightweight version of UNet integrating attention modules and depthwise-separable convolutions, substantially reduces the number of trainable parameters while maintaining high predictive accuracy. SelfAttUNet nie2021self incorporates a self-attention mechanism to highlight key spatial regions for more effective radar-based precipitation prediction. Other studies, such as kaparakis2023wf; reulen2024ga; renault2023sar, have further extended the UNet architecture by combining attention, residual, and generative components, consistently confirming the suitability of encoder–decoder architectures for data-driven nowcasting.
In addition to CNN-based architectures, multi-source data fusion techniques have been increasingly explored to integrate multiple meteorological data sources for enhanced precipitation nowcasting. Traditional models primarily rely on radar reflectivity data; however, recent studies have demonstrated that incorporating additional atmospheric variables, such as wind speed, temperature, and humidity, significantly improves short-term forecasting accuracy. Using multiple data sources, fusion models better capture complex weather dynamics, leading to more reliable nowcast performance. Kaparakis et al. kaparakis2023wf proposed the Weather Fusion UNet (WF-UNet) integrates radar reflectivity and wind maps through a dual-stream UNet framework, demonstrating that multi-variable fusion enhances nowcasting accuracy. leinonen2023thunderstorm introduced a deep learning-based thunderstorm nowcasting model that integrates diverse meteorological data, including weather radar, lightning detection, satellite imagery, numerical weather prediction outputs, and digital elevation models. Large-scale models such as MM-RNN ma2023mm and MetNet-3 andrychowicz2023deep follow this trend by combining radar, satellite, and numerical weather prediction (NWP) data in a single framework to learn relationships across different spatial and temporal scales and to produce more physically consistent forecasts. Furthermore, the Spatiotemporal Feature Fusion Transformer xiong2024spatiotemporal incorporates feature-level fusion mechanisms to enhance precipitation nowcasting performance. Collectively, these works illustrate the growing impact of data fusion techniques in weather nowcasting, highlighting their potential to bridge the gap between numerical weather prediction and deep learning-based methodologies.
3 Methodology
3.1 Dataset
This study utilizes two complementary datasets: a radar-based precipitation dataset and a weather station dataset. Both datasets were obtained from the Royal Netherlands Meteorological Institute (KNMI  [^3] https://www.knmi.nl/home.) and are used together to improve precipitation nowcasting performance.
Precipitation Dataset: The precipitation dataset, originally used by Trebing et al. trebing2021smaat, consists of radar-based rainfall data collected over four years (2016–2019) in the Netherlands. It contains approximately 420,000 rain maps, each captured at 5-minute intervals. The data is sourced from two C-band Doppler weather radar stations located in: De Bilt (52.10°N, 5.18°E, 44 m MSL) and Den Helder (52.96°N, 4.79°E, 51 m MSL). The technical characteristics of these radars including a wavelength of Wavelength (cm), Pulse repetition frequency (Hz), Peak power (kW), Pulse duration ($\mu s$), 3-db beamwidth, Antenna rotation speed of $18\textdegree s^{-1}$, and Number of samples per range ($km^{-1}$). These specifications determine the spatial resolution, sensitivity, and data quality of the radar observations. To generate rain maps, each radar performs four full 360° azimuthal scans at beam elevation angles of 0.3°, 1.1°, 2.0°, and 3.0° precipitation_dataset; trebing2021smaat. The rain maps are rain-gauge adjusted with more details being described in overeem2009derivation. This is the same radar-derived precipitation dataset used in Trebing et al. (2021) and serves as the precipitation input for the models in this study.
Weather Station Dataset: The second dataset consists of weather station data, which provides multivariable atmospheric measurements from 22 weather stations across the Netherlands (see Fig. 1).This dataset contains 8 meteorological variables, selected based on availability, relevance, and consistency. The recorded variables are: Air Temperature, Humidity, Atmospheric Pressure, Average wind Speed, Wind Direction, Standard Deviation Wind Direction, Standard Deviation Wind Direction, Max Wind Speed.
Figure 1: An overview map of all the nodes included in the weather station dataset. The dataset contains a total of 22 weather stations. Some stations share identical coordinates, resulting in 17 nodes in the figure.
Dataset Integration and Processing: To create a unified dataset for model input, the precipitation and weather station datasets undergo several preprocessing steps to ensure consistency and alignment. The eight meteorological variables recorded at each station capture local atmospheric conditions that influence the formation and movement of precipitation. These variables provide complementary information that is not present in the radar sequence alone. For the precipitation maps, a 288×288 pixel cutout is taken from the center of the Netherlands, covering the area with the highest data availability. The selected geographic boundaries are: Longitude: 4.29°E to 6.73°E and Latitude: 51.32°N to 52.81°N. These maps are then resized to 64×64 pixels to reduce memory requirements, as both datasets are incorporated into the models. Finally, the precipitation values are normalized using the maximum precipitation value in the dataset (47.83 mm). The precipitation dataset contains a significant number of radar images with little to no recorded rainfall.
To mitigate the risk of biasing the model toward predicting zero values, precipitation maps where fewer than 50% of the pixels indicate rainfall are excluded. As summarized in Table 1, applying this filtering criterion reduces the full 4-year dataset from 314,940 training samples and 105,003 test samples to 5,734 and 1,557 samples, respectively. Once filtered, the remaining precipitation maps are temporally matched with the corresponding weather station data recorded at the same timestamp. Any data point that does not have a corresponding entry in the other dataset is removed, ensuring that only fully aligned data pairs are retained. To maintain consistency across variables, all weather station measurements are standardized based on their respective means and standard deviations. A challenge in merging the two datasets arises from their differing temporal resolutions. While precipitation maps are recorded at 5-minute intervals, weather station data is available at 10-minute intervals. To resolve this mismatch, the weather station dataset is duplicated to align with the precipitation timestamps, allowing for seamless integration of both sources.
In SmaAt-fUsion, the time series of all meteorological variables across all stations is transformed into a compact 3D tensor, allowing the model to learn how local atmospheric conditions modulate the high-level features at the bottleneck of the SmaAt-UNet. This fusion enriches the latent precipitation representation with complementary physical signals that are not contained in the radar sequence alone. In SmaAt-Krige-GNet, each meteorological variable at each timestep is interpolated using Kriging to create spatially continuous 64 × 64 maps. This process propagates local station measurements across the spatial domain, enabling the model to learn how spatial gradients in temperature, humidity, wind, and pressure relate to evolving precipitation structures. After preprocessing, the final dataset is divided into a training set (2016–2018) and a testing set (2019). During model training, a validation set is created by randomly selecting 10% of the training data.
Figure 2: Overview of the SmaAt-fUsion architecture. Weather station data is integrated into the model by concatenating its learned representation with that of the precipitation maps at the bottleneck of the SmaAt-UNet model. The precipitation stream processes 12 radar maps through the standard SmaAt-UNet encoder–decoder: depthwise-separable convolutions (blue arrows), max-pooling (brown), bilinear upsampling (green), and skip connections (grey). Features progressively compress through the encoder and are then expanded during decoding, forming the main spatial–temporal pathway of the network. In parallel, the station-data stream transforms the 22×8×12 tensor using two 3D depthwise-separable convolutions (pink) and an AdaptiveMaxPool3D layer to produce a compact embedding. This embedding flows into the fusion stage, where it is concatenated with the UNet bottleneck, creating a shared latent representation used for the final precipitation prediction.
Table 1: Size of each train and test set for the 4-year precipitation dataset
Rain Threshold Train Size Test Size Subset $0\%$ (Original) 314940 105003 100% $50\%$ 5734 1557 1.74%
3.2 Proposed models
In the following, we introduce two architectures designed to extend the core SmaAt-UNet model trebing2021smaat to learn from both weather station data and precipitation maps. While both models take 12 radar images (representing one hour of weather radar data) as input and predict the precipitation field 30 minutes ahead, they differ in how the station observations are encoded and integrated. SmaAt-fUsion embeds the weather station variables directly into the UNet bottleneck, allowing meteorological variables to influence high-level latent features. SmaAt-Krige-GNet, in contrast, incorporate station measurements into spatially radar maps using Kriging and processes them through a parallel encoder, enabling multi-scale fusion across the entire UNet architecture.
(a) Overview of the Kriging process. For each variable, a Kriging map is generated for each timestep. Subsequently, 12 timesteps are fed into the SmaAt-Krige-GNet model.
(b) The SmaAt-Krige-GNet model. The upper stream processes Kriging data and mirrors the layers of the bottom stream, which handles precipitation data. At each level, the output of the upper stream is concatenated channel-wise with the corresponding output of the bottom stream.
Figure 3: An illustration of the Kriging process and Overview of the SmaAt-Krige-GNet architecture.
3.3 SmaAt-fUsion model
In SmaAt-fUsion, our aim is to integrate weather station data into the core SmaAt-Unet model trebing2021smaat, which operates on gridded precipitation radar maps. To this end, following the approach of mehrkanoon2019deep, we exploit the spatio-temporal structure of station data by transforming each regressor vector, containing historical weather variable measurements from all stations, into a tensor with dimensions (stations = 22, variables = 8, lags = 12). This tensor is processed through a double 3D depthwise separable convolution block, consisting of a depthwise convolution followed by a pointwise convolution. Compared to classical 3D convolutions, this decomposition significantly reduces the number of parameters while preserving spatial and temporal correlations in the data. The output from this convolutional block is then passed through a 3D Adaptive Max Pooling layer, which reshapes it to match the bottleneck tensor size of the original SmaAt-Unet model. Finally, the pooled representation is concatenated with the bottleneck tensor, as illustrated in Fig. 2. By integrating station data, we enrich the model’s latent space with additional insights from weather station observations. This fusion strengthens the model’s ability to better capture complex spatio-temporal dependencies, ultimately improving its performance in precipitation mapping.
3.4 SmaAt-Krige-GNet model
In this work, we explicitly integrate spatial dependencies between weather stations into the SmaAt-UNet model. To achieve this, we first generate grid-based maps of station variables using Kriging Isaaks1989; Matheron1963. Inspired by reulen2024ga, which introduced SmaAt-GNet as a generator within an adversarial learning framework for precipitation nowcasting, we propose SmaAt-Krige-GNet. This model combines the grid-based maps derived from station data with precipitation maps. The key distinction between SmaAt-GNet and the standard SmaAt-UNet model lies in the addition of an extra encoder stream. This new stream mirrors the encoder layers of SmaAt-UNet and concatenates its representations at each level with those of the original SmaAt-UNet encoder. While the original SmaAt-GNet in reulen2024ga was designed to integrate binary precipitation masks into the SmaAt-UNet model, our proposed SmaAt-Krige-GNet instead utilizes Kriging-based spatial maps. By structuring the model into two independent streams rather than stacking all inputs channel-wise in a single-stream architecture, SmaAt-Krige-GNet can independently learn hidden features from both weather and precipitation data. Since all models process 12 input images, we modify the secondary input of SmaAt-GNet to also accept 12 images. Given that there are 8 weather variables, the resulting input forms a 4D tensor of size $64\times 64\times 8\times 12$, which is then flattened into a $64\times 64\times 96$ layer, as illustrated in Fig. 3(b). To explicitly incorporate spatial data, Kriging is applied to interpolate weather station measurements into 2D geospatial maps, ensuring that each weather variable at each timestep is represented in the same spatial resolution as the precipitation maps.
Table 2: Parameter size of each model
Model Parameters Relative Size
SmaAt-UNet 4.0M 1x
SmaAt-fUsion 5.6M 1.4x
SmaAt-Krige-GNet 11.6M 2.9x
Table 3: Model performance metrics across different precipitation thresholds for the studied dataset.
Threshold Model MSE (pixel) $\downarrow$ F1 Score $\uparrow$ CSI $\uparrow$ HSS $\uparrow$ MCC $\uparrow$
$\geq 0.5\text{ mm/h}$ Persistence 0.065986 0.678436 0.513359 0.241484 0.483398
PySTEPS 0.059155 0.652488 0.484216 0.224309 0.449906
RainyMotion 0.062520 0.678437 0.513359 0.241484 0.483398
SmaAt_UNet 0.043293 0.763476 0.617437 0.299234 0.601438
SmaAt-fUsion 0.035778 0.768003 0.623381 0.297824 0.604511
SmaAt-Krige-GNet 0.035618 0.780774 0.640385 0.315630 0.633025
$\geq 10\text{ mm/h}$ Persistence - 0.117185 0.062239 0.056982 0.113965
PySTEPS - 0.151921 0.082205 0.074494 0.149230
RainyMotion - 0.117211 0.062254 0.056994 0.113989
SmaAt_UNet - 0.144665 0.077972 0.071221 0.148874
SmaAt-fUsion - 0.187187 0.103258 0.092433 0.189564
SmaAt-Krige-GNet - 0.052808 0.027120 0.026092 0.091192
$\geq 20\text{ mm/h}$ Persistence - 0.033757 0.017169 0.016623 0.033265
PySTEPS - 0.050739 0.026030 0.025150 0.050538
RainyMotion - 0.033758 0.017169 0.016623 0.033265
SmaAt_UNet - 0.047089 0.024113 0.023404 0.051688
SmaAt-fUsion - 0.052607 0.027014 0.026190 0.061962
SmaAt-Krige-GNet - 0.029783 0.015117 0.014851 0.054041
Table 4: Ablation study for the eight meteorological variables in SmaAt-Krige-GNet and SmaAt-fUsion at threshold $\geq 0.5\text{ mm/h}$. Removing each variable reveals its relative contribution to precipitation nowcasting performance.
[…]
SmaAt-Krige-GNet None (all variables) 0.7808 0.6404
Humidity 0.7638 0.0170 2.18% 0.6200 0.0204 3.19%
Atmospheric pressure 0.7639 0.0169 2.16% 0.6222 0.0182 2.84%
Wind direction 0.7646 0.0162 2.07% 0.6222 0.0182 2.84%
Avg. wind speed 0.7650 0.0158 2.02% 0.6226 0.0178 2.78%
Max. wind speed 0.7727 0.0081 1.04% 0.6296 0.0108 1.69%
Std. wind speed 0.7801 0.0007 0.09% 0.6403 0.0001 0.02%
Air temperature 0.7808 0.0000 0.00% 0.6404 0.0000 0.00%
Std. wind direction 0.7808 0.0000 0.00% 0.6404 0.0000 0.00%
SmaAt-fUsion None (all variables) 0.7680 0.6231
Wind direction 0.7535 0.0145 1.89% 0.6088 0.0143 2.29%
Atmospheric pressure 0.7539 0.0141 1.84% 0.6080 0.0151 2.42%
Humidity 0.7552 0.0128 1.67% 0.6074 0.0157 2.52%
Avg. wind speed 0.7572 0.0108 1.41% 0.6062 0.0169 2.71%
Max. wind speed 0.7608 0.0072 0.94% 0.6159 0.0072 1.16%
Air temperature 0.7680 0.0000 0.00% 0.6231 0.0000 0.00%
Std. wind direction 0.7680 0.0000 0.00% 0.6231 0.0000 0.00%
Std. wind speed 0.7680 0.0000 0.00% 0.6231 0.0000 0.00%
[…]
4 Experimental setup and evaluation
4.1 Model setup
The models were implemented and trained using PyTorch Lightning. The training process used Adam Optimizer with a learning rate of 0.001. To avoid overfitting, early stopping is applied with a maximum patience of $ES=12$ to monitor the loss of validation. In addition, we apply learning rate scheduling with a patience of $LR=8$ epochs, which reduces the learning rate when no improvement in validation loss was observed.
The models were trained on the dataset for a maximum of 200 epochs, with a batch size of 16. These hyperparameters were chosen based on preliminary experimentation to balance model performance and computational efficiency.
To generate the Kriging maps for SmaAt-Krige-GNet, we use Ordinary Kriging from PyKrige pykrige. A grid is defined with the same resolution as the precipitation maps, specifically $64\times 64$. Subsequently, a variogram is fitted using the spherical model, which has been shown to be suitable for meteorological applications Verworn2011; CSEGRecorderVariograms. To ensure accurate spatial representation, we specify the coordinate type as geographic, allowing the model to directly use the latitude and longitude of each weather station. However, Kriging with PyKrige can be numerically unstable when using its default inverse matrix algorithm. To mitigate this, we set the ‘pseudo_inv’ parameter of Ordinary Kriging to True, which stabilizes computations at the cost of increased processing time pykrige. Once the linear system is solved, a set of weights $\lambda_{i}$ is obtained, which are then applied to interpolate the Kriging value at each coordinate within the grid. This process generates spatially continuous weather variable maps that align with the resolution of the precipitation data.
Figure 4: An example of Kriging map input of the eight meteorological variables at a selected timestamp. To improve visibility, Atmospheric pressure uses a percentile-based color scale due to its narrow dynamic range, while the other variables share a common normalization (2-2 to 2).
4.2 Baselines
To comprehensively evaluate the performance of the proposed models, we compared them against four representative baselines both deep learning and optical flow approaches. The Persistence model simply assumes that future precipitation will remain the same as the latest observation. SmaAt-UNet trebing2021smaat represents a compact attention-based UNet widely adopted in radar nowcasting. In addition, two optical flow methods were included: PySTEPS pulkkinen2019pysteps which performs stochastic extrapolation using motion fields, and RainyMotion ayzel2019optical, a CNN-based flow estimation model. Each nowcasting case consists of 12 input radar frames (one hour at 5-minute intervals) and a single prediction at +30 minutes, consistent with prior UNet-based nowcasting studies. Thus, the temporal frequency of evaluation is also 5 minutes, matching the native radar observation cycle, ensuring a fair comparison across architectures.
4.3 Evaluation
The loss function utilized in this study is the Mean Squared Error (MSE) hodson2021mean, which quantifies the difference between the predicted and ground truth images. In addition to MSE, and in line with previous works reulen2024ga; trebing2021smaat; stanczyk2021deep; van2023improving, we employ a set of widely used binary metrics for precipitation nowcasting, including the F1-Score naidu2023review, Critical Success Index (CSI) jolliffe2003, Heidke Skill Score (HSS) hogan2012, and Matthews Correlation Coefficient (MCC) mcc, to assess the performance of the proposed models. Since precipitation maps contain continuous values, they must be thresholded to convert them into binary masks before applying these metrics.
In this study, we evaluate precipitation at three different thresholds: 0.5 mm/h, 10 mm/h, and 20 mm/h. Applying a threshold to the predicted and target images produces a binary mask, where each pixel is assigned 1 if its value exceeds the threshold (rain) and 0 otherwise (no rain). From this, one can calculate the true positives (TP) (prediction=1, target=1), false positives (FP) (prediction=1, target=0), true negatives (TN) (prediction=0, target=0) and false negatives (FN) (prediction=0, target=1). To ensure statistical robustness, these contingency counts are first accumulated over all 1,557 nowcasting cases from the full 2019 evaluation period, rather than being computed on a per-sample basis. The final CSI, HSS, and MCC values are then obtained by applying the metric formulas to these aggregated pixel-wise totals, resulting in stable estimates that reflect performance across diverse precipitation events. These metrics also guarantees that the reported scores represent the entire annual distribution of rainfall. In the 2019 test set, this distribution is reflected by 263 winter cases (16.9%), 504 spring cases (32.4%), 342 summer cases (22.0%), and 448 autumn cases (28.8%), demonstrating that the evaluation covers all four seasons and a diverse range of meteorological conditions.
Using thresholds, the models’ performance can be assessed across varying precipitation intensities. The 0.5 mm/h threshold is commonly used as it effectively captures light rain tan2018evaluation. Additionally, 10 mm/h and 20 mm/h thresholds are tested to evaluate the models’ ability to predict moderate to extreme precipitation events. The binary metrics CSI, HSS and MCC are computed as follows:
CSI $=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}},$ (1)
HSS $=\frac{2\cdot(\text{TP}\cdot\text{TN}-\text{FP}\cdot\text{FN})}{(\text{TP}+\text{FN})\cdot(\text{FN}+\text{TN})+(\text{TP}+\text{FP})\cdot(\text{FP}+\text{TN})},$
MCC $=\frac{\text{TP}\cdot\text{TN}-\text{FP}\cdot\text{FN}}{\sqrt{(\text{TP}+\text{FP})(\text{TP}+\text{FN})(\text{TN}+\text{FP})(\text{TN}+\text{FN})}}.$
The CSI evaluates the overlap between predicted and observed precipitation events jolliffe2003. It accounts for false positives and missed events, providing a robust measure for precipitation predictions hogan2012. The HSS compares the model predictions to a random baseline hogan2012. HSS highlights the skill of the model beyond chance-level prediction wilks2011. Unlike accuracy, which can be misleading in imbalanced datasets, the MCC score provides a more reliable measure by considering both classes (positive and negative). The MCC ranges from -1 (perfect inverse correlation) to +1 (perfect correlation), with 0 indicating no better than random predictions. This metric is particularly useful when the data is highly imbalanced, as it accounts for both the correct and incorrect predictions across all categories mcc. In the case of precipitation nowcasting, the present imbalance is the high amount of negative precipitation events compared to positive precipitation events.
Figure 5: Examples precipitation input of the dataset on 2019-03-10, 2019-06-15, 2019-09-29, 2019-12-14.
Figure 6: Example outputs of each model, using the input seen in Fig. 5
5 Results and Discussion
5.1 Model performance
After training, models were selected based on the lowest validation loss to ensure optimal generalization. The obtained results are presented in Table 3. These results were computed after denormalizing the data, following the approach used in similar studies reulen2024ga; trebing2021smaat.
The results in Table 3 show that the proposed SmaAt-fUsion model outperforms the classical SmaAt-UNet across all tested thresholds. This suggests that incorporating station data effectively guides the SmaAt-UNet core model in better learning atmospheric dynamics. Furthermore, the SmaAt-fUsion model achieves superior performance compared to SmaAt-Krige-GNet at higher rainfall thresholds of 10 mm/h and 20 mm/h. However, for low-intensity rainfall, SmaAt-Krige-GNet performs best among the examined models. This indicates that integrating additional weather variables at multiple levels of the SmaAt-UNet encoder enables the model to extract more detailed information from the augmented weather station data, compared to the SmaAt-fUsion model.
At higher rainfall thresholds (10 mm/h and 20 mm/h), SmaAt-Krige-GNet underperforms relative to both SmaAt-fUsion and the core SmaAt-UNet in terms of binary evaluation metrics. This performance decrease is likely due to the model’s tendency to over-smooth predictions, primarily a consequence of incorporating Kriging-generated maps. As shown in Fig. 4, these maps inherently introduce smoothness. Although Kriging provides spatial continuity required by convolutional architectures, the limited number of station observations reduces its ability to reconstruct fine-scale atmospheric structures. This smoothing effect is especially noticeable in variables such as atmospheric pressure and wind-related fields, where local gradients are reduced. Consequently, SmaAt-Krige-GNet benefits from these maps in low-intensity rainfall scenarios, where larger-scale patterns dominate but struggle with sharp precipitation peaks, which require higher spatial resolution. Increasing the density of weather stations would directly improve the fidelity of the interpolated fields. A potential solution to this limitation would be a significant increase in the number of weather stations, which could improve the resolution and accuracy of the interpolated Kriging maps.
Fig. 5 and Fig. 6 further clarifies the strengths and remaining limitations of the proposed architectures. Fig. 5 illustrates four representative precipitation maps across different seasons. These inputs contains both stratiform and convective rainfall, providing a challenging test for models relying solely on reflectivity patterns. 2019-03-10, 2019-09-24, and 2019-12-14 correspond to stratiform precipitation events associated with large-scale frontal systems, characterized by broad rainfall distributions. In contrast, 2019-06-15 represents a convective summer case with sharp spatial gradients. Including both regimes enables a comprehensive evaluation of the models’ ability to handle diverse rainfall patterns.
In Fig. 6, most of models correctly capture the general location and shape of the main precipitation areas. However, SmaAt-fUsion and SmaAt-Krige-GNet produce blurred predictions, particularly along convective boundaries. This is most apparent in the summer case (2019-06-15), where sharp convective streaks are substantially smoothed out. By contrast, PySTEPS and RainyMotion do not suffer from such over-smoothing. Their outputs preserve sharper gradients due to the nature of optical-flow extrapolation. Nevertheless, their predictive skill is limited. Both methods show displacement errors, misplacing the precipitation band in several cases, and in some situations their performance is nearly indistinguishable with the Persistence baseline. These results highlight the trade-off between spatial sharpness and predictive accuracy among different nowcasting approaches.
Recent generative nowcasting models such as DGMR ravuri2021skilful, DiffCast yu2024diffcast, and NowcastNet Zhang2023 produce sharper and more realistic predictions, particularly at longer lead times. These models also achieve strong performance without requiring additional meteorological inputs. In contrast, our approach focuses on multi-data integration by incorporating weather-station variables, which strengthens the physical basis of the predictions and improves performance in low-intensity scenarios. Combining generative architectures with our fusion strategies is a promising direction, as generative models could reduce smoothing and peak underestimation at longer lead times while benefiting from the added atmospheric information provided by multi-modal datasets.
5.2 Variable Analysis
To quantify the individual contribution of each meteorological variable, we conduct a study in which each of the eight weather-station variables is removed while all other inputs remain unchanged. The ablation results in Table 4 provide further insight into how individual meteorological variables influence the performance of the two proposed models. Across both architectures, removing a single variable results in a decline in predictive skill, indicating that the station-derived features meaningfully complement the radar-based inputs. The magnitude of this decline reveals clear differences in the roles played by specific atmospheric variables.
For SmaAt-Krige-GNet which incorporates spatially Kriging interpolated weather shows the strongest sensitivity to humidity, atmospheric pressure, wind direction, and average wind speed. Removing any of these variables produces a noticeable degradation ($\Delta F1\approx 0.016–0.017$), corresponding to a contribution of roughly 2–3%. In contrast, variables that represent short-term fluctuations such as the standard deviations of wind speed and direction have less contribution to the model. The SmaAt-fUsion model shows a similar ranking of variable importance, though with generally smaller degradation when variables are removed. The reason is SmaAt-fUsion integrates station information only at the bottleneck rather than at multiple spatial scales. Nevertheless, the same four variables wind direction, atmospheric pressure, humidity, and average wind speed—remain the most influential. Removing these variables reduces F1 by approximately 0.011–0.015, corresponding to 1.4–1.9% contribution. The consistency of variable importance across both architectures indicates that our multi-source data fusion strategy effectively exploits the complementary strengths of weather-station variables and radar information.
6 Conclusion
This study introduced two novel deep learning models, SmaAt-fUsion and SmaAt-Krige-GNet, designed to improve precipitation nowcasting by integrating meteorological variables from weather stations alongside radar data. The experimental results demonstrate that both models have distinct advantages: SmaAt-Krige-GNet excels in low-intensity rainfall predictions by incorporating weather station data through Kriging-based spatial interpolation, while SmaAt-fUsion outperforms in high-intensity events by directly integrating station data into the network’s bottleneck. These findings highlight the potential of leveraging both discrete weather station data and spatially interpolated variables to improve deep learning-based precipitation forecasting. Future work includes extending the proposed fusion strategies to alternative architectures, additional multi-source meteorological datasets, and longer lead times to further assess their generality and robustness across diverse nowcasting settings.